the perimeter of a rectangle is 500 yards. What are the dimensions of the rectangle if the length is 50 yards more than the width?

Answer by checkley79(3341)

P=2L+2W
L=W+50
500=2(W+50)+2W
500=2W+100+2W
500=4W+100
4W=500-100
4W=400
W=400/4
W=100 ANS.
L=100+50=150 ANS.
PROOF:
500=2*150+2*100
500=300+200
500=500